A number is divided by 3. 5 is added to the quotient. When the sum is multiplied by 2, the result is 60. What is the number?
Answer by richwmiller(17219)

(n/3+5)*2=60
n/3+5=30
n/3=25
n=75
check
75/3=25
25+5=30
2*30=60
60=60
ok